The sum of two numbers is 50. One number is 5 less than 4 times the other number. Find the numbers.
"I get stuck right at the beginning, because I'm not sure whether I use 2 variableas or just one! I know how to set up the first sentence, but the second one confuses me. "

Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

The sum of two numbers is 50. One number is 5 less than 4 times the other number. Find the numbers.
------
Equations:
x + y = 50
x = 4y-5
-----
Substitute for "x" and solve for "y":
4y-5+y = 50
5y = 55
y = 11
----
Solve for "x":
x = 4y-5
x = 4*11-5
x = 39
=============
Cheers,
Stan H.

Answer by ewatrrr(24785)

 
Hi
The sum of two numbers is 50. x+y = 50 and y = (50-x)
regardless, starting with one or 2, we would using one to complete
Let x and (50-x) represent the two numbers.
x = 4(50-x)-5
5x = 195
x = 39 the numbers are 39 and 11
39 = 4*11 - 5 = 44-5